OpenStudy (anonymous):

Darlene has a balance of $3980 on a credit card with an APR of 22.8%. Paying off her balance in which of these lengths of time will result in her paying the least amount of interest? A. 9 months B. 6 months C. 5 months D. 8 months @caozeyuan @Directrix @e.cociuba

Directrix (directrix):

I would like to say 5 months because the quicker these card balances are paid off, the less the interest. But, there are all sorts of tricks (my opinion) that these companies play. Let me go over to the online credit card site and see what I can find there. Hold on.

Directrix (directrix):

I'm over here working in case you are wondering: http://www.bankrate.com/calculators/credit-cards/credit-card-payoff-calculator.aspx

Directrix (directrix):

Wow. This work assumes that no additional charges will be made to the account during the stated time period. I found the monthly payments for the given months at the website posted above. The monthly payment included the interest but did not state the amount. So, I multiplied the monthly payment for each of the time periods and from that subtracted the amount owed to get the interest. As we suspected all along, the shortest payoff period results in the least amount of interest. I thought you might want to see just how much interest for each payoff period. 485.29 * 9 -3980= 387.61 841.94 * 5 -3980 = 229.70 ---> Five months 708.14 * 6 -3980= 268.84 540.97 * 8 -3980= 347.76

Directrix (directrix):

@divagirl421